OpenStudy (anonymous):

What impulse is exerted in each of the following cases? The Earth pulling down on a 12kg rock during the 3.0 seconds it takes to fall down from a cliff A billiard ball of mass 200 g rolls towards the right hand cushion of a billiard table at 2.0m/s and rebounds straight back at 2.0m/s. (a) What is its change in momentum as a result of hitting the cushion. (b) what impulse is given to the ball by the cushion Please help, will give medals out but I'm so confused by this completely.

OpenStudy (anonymous):

Impulse is the same as momentum; impulse or pulse is a sudden increaseof a quantity followed by a decrease then it comes another increase, occuring alternatively, just like the heart beat reading in that machine.

OpenStudy (anonymous):

So pulse is Ft, where F is force and t time. Or mv, where m is mass and v is velocity. We can use Ft, because we can find F and we have t. F is mg or 12*9.81. And t is 3seconds. Hence, we ultimately have Ft or 12*9.81*3=?

OpenStudy (anonymous):

Impulse is the change of momentum or Force over a period of time.

OpenStudy (anonymous):

If a rock of 12 kg falls from a cliff for 3s, the impulse is 12*9.8*3=352.8

OpenStudy (anonymous):

impulse is like 'change in force in a short period of time' so, for an impulse like take 'Im' = F * t (force * time) and can also be written as im = m * v (mass * Velocity) which is also momentum = M therefore: a) momentum = m * v ( here v is not 2.0 m/sec. since, it is difference in velocity, i.e., v1-v2 = 2.0 - 2.0 = 0) so, momentum = 200 * 0 = 0 hence, no change in momentum... and for the rock question... the first guy's answer deserve a medal... Keep Asking & Answering...

OpenStudy (anonymous):

Thanks.